solution z is prepared by mixing 0.15dm3 of solution X and 0.35 dm3 of solution Y. Solution X contains 0.12 mole of ammonia, NH3 while solution Y contains 0.09 moles of ammonium chloride, NH4CL. Determine the pH of solution Z after the addition of 0.01 mole of hydrochloric acid, HCl.( Given the pKb of NH3 is 4.75)



1
Expert's answer
2021-02-05T04:01:59-0500

NH3 + NH4Cl - buffer solution

NH3 + HCl = NH4Cl

0.12 mole 0.01 mole 0.01 mole

HCl - limiting reagent

n(NH3) = 0.12 mole - 0.01 mole = 0.11 mole

n(NH4Cl) = 0.09 mole + 0.01 mole = 0.10 mole

V(solution) = 0.15 L + 0.35 L = 0.5 L

C(NH3) = n(NH3) / V(solution) = 0.11 mole / 0.5 L = 0.22 M

C(NH4Cl) = n(NH4Cl) / V(solution) = 0.10 mole / 0.5 L = 0.20 M

pOH = pKb + log (molarity of salt / malarity of base) = 4.75 + log (0.20 / 0.22) = 4.71

pH = 14 - 4.71 = 9.29